# T-FAKE: Synthesizing Thermal Images for Facial Landmarking

### v2

@@ 5.4 Comparison for Thermalization
Our approach aims to overcome the limitations of thermal datasets to fixed camera angles, controlled lighting, and stable temperatures. However, this data limitation hinders the evaluation of ‘in-the-wild’ images. Therefore, we report three different metrics. Firstly, we report the average Fréchet Inception Distance (FID) [21] between all available thermal SEJONG images and five different equally-sized thermalized FAKE subsets. The FID measures image translation without ground truth image pairs.Here, it compares the perceptual similarity of the thermal SEJONG dataset and a dataset of thermalized FAKE images. Additionally, we report the mean-squared error (MSE) for two paired thermal datasets (with filtered background) to evaluate actual out-of-distribution prediction quality. Here, we use TUFTS [48] and 1000 randomly selected images from DRIVE-IN [13]. TUFTS has been recorded under conditions similar to SEJONG and DRIVE-IN features side views from people in cars. Table 4 compares our final model ($\lambda_{R}=1$, $\lambda_{W}=0.01C$) to two supervised models trained with SEJONG only, a baseline implementation ($\lambda_{R}=0$, $\lambda_{W}=0$) and a Pix2Pix model [26], and three unsupervised models trained with SEJONG and FAKE, namely CycleGAN [66], CUT [49] and QS-Attn [23]. The MSE shows that our model outperforms all other models for temperature prediction and displays domain adaptation for the out-of-distribution DRIVE-IN data recorded outside a lab. Moreover, the resulting T-FAKE images display better perceptual similarity than images produced by all other methods except for QS-Attn. However, QS-Attn fails at temperature prediction, especially for the non-frontal DRIVE-IN images.
@@ 5.5 Thermalization Ablation Study
As an additional ablation of our regularizers, Table 4 presents all the previously introduced metrics for different choices of $\lambda_{W}$ and $\lambda_{T}$. We see that the segmentation-based regularizer leads to a large boost in perceptual similarity and temperature prediction power. The patch-based regularizer leads to a smaller boost. The combination of both regularizers always gives the best or second-best result. This confirms the effectiveness of our RGB2Thermal loss.
@@ 6 Conclusion
To overcome data limitations in thermal landmarking, we established a semi-supervised RGB2Thermal image translation method. Using patch-based and segmentation-based regularization, we ensured the generalization of our thermalizer to the synthetic FAKE dataset. We employed this model to generate the up-to-date largest thermal landmarking dataset, the T-FAKE dataset. By training a probabilistic landmarker with the T-FAKE dataset, we established the up-to-date first dense thermal landmarker and a state-of-the-art RGB+Thermal sparse landmarker. We tested these landmarkers with strict benchmarking by including established thermal and RGB methods and using label adaptation for each benchmarked method. Our work shows that synthetic data is sufficient for the training of thermal landmarkers and can be done on multiple imaging modalities.
@@ A.2 Thermalization Implementation Details
For the patch-based regularizers, we use a patch size of $8$. We disregard the background by excluding synthetic patches based on the ground truth background segmentation and completely black real patches. For our multi-scale approach, we sum the regularizer over $5$ scales with a downsampling factor of $0.5$. Furthermore, facial temperatures depend on the surrounding temperature. Thus, we train a ‘warm’ and a ‘cold’ model with different reference temperature values for the segmentation-based regularizers. For data augmentation, we use the same random rotations and cropping for the natural RGB and thermal images. Moreover, we apply random color changes, blurring, and shadow augmentations [43] exclusively to the natural RGB images. Next, we apply random rotations and cropping for the synthetic RGB images. Here, we also fill holes in the original ‘glasses’ segmentation masks showing outlines of the frame only to highlight transparent, but heat-blocking glass or plastic. Lastly, we replace the original background with a black background based on the known semantic segmentation. We use a U-Net $T_{\theta}$ with a Resnet34 encoder pre-trained on ImageNet [10] and train it for 10 FAKE epochs with SEJONG batches of size $64$ and FAKE batches of size $64$. This corresponds to approximately 100 SEJONG epochs. Further, we use an Adam optimizer with an initial learning rate of 0.001 which we reduce to 0.0001 after 4 epochs. Based on a random split, we use 80% of the SEJONG data for training and all available FAKE data. Also, we use the geomloss [12] implementation for the Wasserstein patch loss with $\lambda_{E}=0.1^{6}$. We normalize the squared error loss $\|\cdot-\cdot\|_{2}^{2}$ by dividing with the image dimension, here $256^{2}$. Based on the $5$ scales and patch dimension $8^{2}$, we set $\lambda_{W}=0.01C$ for our final model with the normalization constant $C=(5\cdot 8^{2})^{-1}$. We set $\lambda_{R}=1$. Note that the choice of $\lambda_{R}=1$ and $C$ are motivated to control the value range. The (normalized) MSE data fidelity term $\|\cdot-\cdot\|_{2}^{2}$ which we evaluate on normalized SEJONG images takes values in the range $[0,1]$ for arbitrary images with pixel range $[0,1]$ due to dimensional normalization. To have a similar value range for the evaluated FAKE images, the patch-based regularizer $W$ takes values in $[0,1]$ for $\lambda_{W}=C$ with arbitrary normalized images. Moreover, the segmentation-based regularizer $R$ also takes values in $[0,1]$ for such images given arbitrary normalized reference temperature values in $[0,1]$ for $\lambda_{R}=1$.
@@ A.3 SEJONG Dataset
The SEJONG dataset illustrates the impact of various disguises, including glasses, wigs, and fake beards. Each subject in the dataset is presented with different disguises. As a result, it includes a lot of clothing and hairstyle variations. Most participants have a Southeast Asian or Central Asian ethnic background, but people of other ethnicities are included too. The number of participants identifying as male or female is balanced. This makes it an attractive dataset candidate for thermalization training that is supposed to generalize to a large variety of subjects. Due to reasons of data privacy, we refer to the original publication [7] and have to abstain from showing randomly selected SEJONG images.
@@ A.6 Extended Thermalization Ablation Study
NME W/H $\downarrow$ FAKE[61, 62] FAKE 0.1055 0.2675 0.1241 0.2534 0.1887
FAKE* FAKE 0.0933 0.2682 0.1312 0.2348 0.1824
Thermal Only FAKE + T-FAKE 0.0832 0.1334 0.0677 0.1503 0.1090
RGB + Thermal FAKE + T-FAKE 0.0740 0.1346 0.0684 0.1420 0.1051
Table 8: Ablation results on CHARLOTTE dataset splits. Pre-processing with the pre-processing stack for RGB landmarkers is indicated by *. The confidence threshold has been set to infinity. RGB + Thermal and Thermal Only models have been finetuned from the FAKE models.
@@ Appendix C Large-Scale Visualization
For a large number of T-FAKE samples, we refer to Figures 16 and 17. Here, we simply use the first 128 images based on the numerical naming convention of the original FAKE dataset.
@@ Appendix E Limitations
Thermalization. This work depends on the thermalization of the final renders in the FAKE dataset. The dataset contains very difficult lighting conditions and scene compositions that make it powerful to train landmarkers but also made the thermalization particularly challenging and could only be solved with advanced domain-adaptive semi-supervised regularization approaches. Despite a good perceptual result of the faces, some T-FAKE images can contain minor artifacts on clothing and on the background (e.g. see Figure 12, bottom second from right). Nevertheless, these artifacts remain limited. Moreover, background artifacts can easily be removed by choosing a suitable background based on the ground truth segmentation as implemented during our landmarker training, see Fig. 18. We merely used an MSE loss for our supervised training. Including an adversarial [26, 64] or a perceptual [27, 53] loss into our model should lead to additional perceptual improvement.
Dense Landmarks. In this work, we relied on the ground truth 70-point landmarks of the FAKE dataset [61] and dense Mediapipe [40] annotations. Training with the original 320-point and 702-point landmarks could further boost accuracy and lead to an even denser landmarker. However, these landmarks are not publicly available. Furthermore, we only evaluate a mobilenet backbone for landmark detection without face tracking. Better performance for difficult poses and face variability could be achieved with denser models [62] and spatial normalization of face positions during training.
@@
Appendix F Societal Impact Statement
An increasing number of risks are associated with collecting and using biometric information, such as (thermal) facial images. As for collecting such information in general, biometric datasets contain information that could be misused. Biometric applications, such as facial recognition or landmarking technology, may have internal biases and may display higher error rates for certain populations than for others in terms of, e.g., ethnicity, gender, or age. The use of synthetic data can be very helpful for such challenges as it allows for explicit control of dataset biases and a decreased need for open-source biometric personal information. Nevertheless, facial recognition technologies can be misused for surveillance purposes and criminal activities. We hope that our research is not employed with malicious intent. Therefore, we strongly encourage the legislative regularization of face recognition and face analysis applications.
Nevertheless, we want to point out the advantages of facial analysis in the thermal domain since this domain is nearly invariant to lighting conditions and inherently difficult to falsify. Highly accurate thermal landmarks can offer security and authentication benefits to society. The computation of reliable thermal landmarks is a key technology for medical screening of various diseases and remote assessment of the physiological and psychophysiological state of humans [13]. Contactless screening for infectious diseases has the potential to reduce environmental impact as well as to increase security through upscaled screening in pandemic or endemic scenarios. This, however, requires a detailed understanding of the limitations of such methods and can otherwise result in a false sense of security or discrimination against false positive users and dense landmarks offer a way to better interpret imaged temperatures from the face and pinpoint contributions from distinct facial areas.
The availability of dense thermal imaging to the biomedical research communities promises new applications and new scientific insights from highly accurate temperature analysis of the face. Traditionally, studies involving thermal imaging of faces are analyzed within a semi-automated pipeline which would require manual region-of-interest definitions [25] that limit the application of the results in out-of-lab environments.

### v3

@@ 5.4 Comparison for Thermalization
Our approach aims to overcome the limitations of thermal datasets to fixed camera angles, controlled lighting, and stable temperatures. However, this data limitation hinders the evaluation of ‘in-the-wild’ images. Therefore, we report three different metrics. Firstly, we report the average Fréchet Inception Distance (FID) [20] between all available thermal SEJONG images and five different equally-sized thermalized FAKE subsets. The FID measures perceptual quality without ground truth image pairs. Here, it compares the perceptual similarity of the thermal SEJONG dataset and thermalized FAKE datasets. Also, we report the mean-squared error (MSE) for two paired thermal datasets (with filtered background) to evaluate actual out-of-distribution temperature prediction. Here, we use TUFTS [47] and 1000 randomly selected images from DRIVE-IN [13]. TUFTS features frontal lab-conditon images and DRIVE-IN contains side views from people in cars. Table 3 compares our final model ($\lambda_{R}=1$, $\lambda_{W}=0.01C$) to two supervised models trained with SEJONG only, a baseline implementation ($\lambda_{R}=0$, $\lambda_{W}=0$) and a Pix2Pix model [25], and three unsupervised models trained with SEJONG and FAKE, namely CycleGAN [65], CUT [48] and QS-Attn [22]. Our model outperforms all other models for temperature prediction and displays domain adaptation for the out-of-distribution DRIVE-IN data recorded outside a lab. Moreover, the resulting T-FAKE images display better perceptual similarity than all other methods except for QS-Attn. However, QS-Attn fails at temperature prediction, especially for the non-frontal DRIVE-IN images.
@@ 5.5 Thermalization Ablation Study
As an additional ablation of our regularizers, Table 3 presents FID and MSE for varied $\lambda_{W}$ and $\lambda_{T}$. The segmentation-based regularizer leads to a large boost in perceptual similarity and temperature prediction power. The patch-based regularizer leads to a smaller boost. Combining both regularizers always gives the best or second-best result confirming the effectiveness of our RGB2Thermal loss.
@@ 6 Conclusion
Overcoming thermal data limitations, we established semi-supervised RGB2Thermal image translation. Using patch-based and segmentation-based regularization, we ensured thermalizer generalization to the synthetic FAKE dataset. This model generated the up-to-date largest thermal landmarking and segmentation dataset, the T-FAKE dataset. Using T-FAKE, we established the up-to-date first dense thermal landmarker and a state-of-the-art RGB+Thermal sparse landmarker. We performed strict landmarker benchmarking based on established thermal and RGB methods. Moreover, the underlying RGB2Thermal model displays excellent perceptual and temperature prediction performance.
@@ 7 Acknowledgements
Supplementary Material
In the supplementary material, we give additional information for our method. In Section A we provide more details on the thermalization including implementation details and an extended ablation study. In Section B, we add details on landmarker and label adaptation implementation. In Sections C-14 we discuss limitations and provide additional result images of the datasets.
@@ A.2 Thermalization Implementation Details
For the patch-based regularizers, we use random batches with patch size of $8$. We disregard the background by excluding synthetic patches based on the ground truth background segmentation and completely black real patches. For our multi-scale approach, we sum the regularizer over $5$ scales with a downsampling factor of $0.5$. Again, facial temperatures depend on the surrounding temperature. Thus, we train a ‘warm’ and a ‘cold’ model with different reference temperature values for the segmentation-based regularizers. For data augmentation, we use the same random rotations and cropping for the natural RGB and thermal images. Moreover, we apply random color changes, blurring, and shadow augmentations [42] exclusively to the natural RGB images. Next, we apply random rotations and cropping for the synthetic RGB images. Here, we also fill holes in the original ‘glasses’ segmentation masks showing outlines of the frame only to highlight transparent, but heat-blocking glass or plastic. Lastly, we replace the original background with a black background based on the known semantic segmentation. We use a U-Net $T_{\theta}$ with a Resnet34 encoder pre-trained on ImageNet [10] and train it for 10 FAKE epochs with SEJONG batches of size $64$ and FAKE batches of size $64$. This corresponds to approximately 100 SEJONG epochs. Further, we use an Adam optimizer with an initial learning rate of 0.001 which we reduce to 0.0001 after 4 epochs. Based on a random split, we use 80% of the SEJONG data for training and all available FAKE data. Also, we use the geomloss [12] implementation for the Wasserstein patch loss with $\lambda_{E}=0.1^{6}$. We normalize the squared error loss $\|\cdot-\cdot\|_{2}^{2}$ by dividing with the image dimension, here $256^{2}$. Based on the $5$ scales and patch dimension $8^{2}$, we set $\lambda_{W}=0.01C$ for our final model with the normalization constant $C=(5\cdot 8^{2})^{-1}$. We set $\lambda_{R}=1$. Note that the choice of $\lambda_{R}=1$ and $C$ are motivated to control the value range. The (normalized) MSE data fidelity term $\|\cdot-\cdot\|_{2}^{2}$ which we evaluate on normalized SEJONG images takes values in the range $[0,1]$ for arbitrary images with pixel range $[0,1]$ due to dimensional normalization. To have a similar value range for the evaluated FAKE images, the patch-based regularizer $W$ takes values in $[0,1]$ for $\lambda_{W}=C$ with arbitrary normalized images. Moreover, the segmentation-based regularizer $R$ also takes values in $[0,1]$ for such images given arbitrary normalized reference temperature values in $[0,1]$ for $\lambda_{R}=1$.
@@ A.3 SEJONG Dataset
The SEJONG dataset illustrates the impact of various disguises, including glasses, wigs, and fake beards. Each subject in the dataset is presented with different disguises. As a result, it includes a lot of clothing and hairstyle variations. Most participants have a Southeast Asian or Central Asian ethnic background, but people of other ethnicities are included too. The number of participants identifying as male or female is balanced. This makes it an attractive dataset candidate for thermalization training that is supposed to generalize to a large variety of subjects. Due to reasons of data privacy, we refer to the original publication [7] and have to abstain from showing additional SEJONG images.
@@ A.6 Extended Thermalization Ablation Study
NME W/H $\downarrow$ GLL + RW ($\bar{\sigma}<\infty$) [60, 61] FAKE 0.1055 0.2675 0.1241 0.2534 0.1887
GLL + RW* ($\bar{\sigma}<\infty$) FAKE 0.0933 0.2682 0.1312 0.2348 0.1824
GLL + RW ($\bar{\sigma}<\infty$) T-FAKE 0.0832 0.1334 0.0677 0.1503 0.1090
GLL + RW ($\bar{\sigma}<\infty$) FAKE + T-FAKE 0.0740 0.1346 0.0684 0.1420 0.1051
Table 8: Ablation results on CHARLOTTE dataset splits. Pre-processing with the pre-processing stack for RGB landmarkers is indicated by *. The confidence threshold has been set to infinity. RGB + Thermal (FAKE + T-FAKE) and Thermal Only (T-FAKE) models have been finetuned from the FAKE models.
@@ Appendix C Large-Scale Visualization
For a large number of T-FAKE samples, we refer to Figures 15 and 16. Here, we simply use the first 128 images based on the numerical naming convention of the original FAKE dataset.
@@ Appendix E Limitations
Thermalization. This work depends on the thermalization of the final renders in the FAKE dataset. The dataset contains very difficult lighting conditions and scene compositions that make it powerful to train landmarkers but also made the thermalization particularly challenging and could only be solved with advanced domain-adaptive semi-supervised regularization approaches. Despite a good perceptual result of the faces, some T-FAKE images can contain minor artifacts on clothing and on the background (e.g. see Figure 11, bottom second from right). Nevertheless, these artifacts remain limited. Moreover, background artifacts can easily be removed by choosing a suitable background based on the ground truth segmentation as implemented during our landmarker training, see Fig. 17. We merely used an MSE loss for our supervised training. Including an adversarial [25, 63] or a perceptual [26, 52] loss into our model might lead to perceptual improvement.
Dense Landmarks. In this work, we relied on the ground truth 70-point landmarks of the FAKE dataset [60] and dense Mediapipe [39] annotations. Training with the original 320- and 702-point landmarks could further boost accuracy and lead to an even denser landmarker. However, these landmarks are not publicly available. Furthermore, we only evaluate a mobilenet backbone for landmark detection without face tracking. Better performance for difficult poses and face variability could be achieved with denser models [61] and spatial normalization of face positions during training.